Mr Lebelo's husband is 45 years old and he has a healthy body weight and no health problems. Mrs Lebelo is concerned that her husband is not eating healthily and that this might affect his weight in future. He consumes 210g of carbohydrates, 210g of fat and 240g of protein per day. Calculate the percentage of energy each of the energy yielding nutrients contributes to the energy intake per day.


1
Expert's answer
2021-06-02T00:50:03-0400

 Known

Energy per gram

·        Fat = 9 kcal/g (Cal/g)

·        Carbohydrate = 4 kcal/g (Cal/g)

·        Protein = 4 kcal/g (Cal/g)

Grams

·        Fat = 210 g

·        Carbohydrate = 210 g

·        Protein = 240 g

 

·         Energy from Fat: 

·        210 g x 9 kcal/g = 1890 kcal = 1890 Cal

·        Convert to kJ

·        1890 Cal x 4.184kJ/Cal = 7907.76 kJ= 7908kJ

 

·        Energy from Carbohydrate:

·        210 g x 4 kcal/g = 840 kcal = 840 Cal

·        Convert to kJ

·        840 Cal x 4.184kJ/Cal = 3514.56 kJ = 3515 kJ

 

·        Energy from Protein:

·        240 g x 4 kcal/g = 960 kcal = 960 Cal

·        Convert to kJ

·        960 Cal x 4.184kJ/Cal = 4016.64 kJ = 4017 kJ

Total energy consumed per day=7908+35145+4017

                                                       =15440 KJ

% of fat= 7908/15440(100)

            =51.2%

% of carbohydrates= 3515/15440(100)

                              = 22.8%

% of proteins= 4017/15440(100)

                     =26%